One possible worry about electric cars is that they're currently powered by dirty electricity. But proponents, as quoted in this story on the issue, claim that they're still cleaner than traditional cars even if their electricity comes from fossil fuels because of how efficient large-scale electricity production is and how efficient electric motors are. Let's investigate for ourselves.

1.) Let's look at the worst-case scenario, an electric car powered entirely by electricity from coal. Assume the car is a Nissan Leaf. How many kWh does it take to go 1 km, under "average" driving conditions? Try looking online to find the information that you need for this answer. We'll accept a range of answers for this.

2.) How many kg of CO2e emissions would this lead to? Use the median estimate from the figures at the US National Renewable Energy Laboratory for the emissions intensity of coal electricity.

3.) What are the per-km emissions from a standard American car? Assume a fuel economy of 9.5L/100km, with gasoline that has a lifecycle impact of 3.2 kg CO2e/liter.
Answer #1

Answer 1.)

As per online information available Nissan Leaf car can go for 3 Miles/kWh.

As 1 Mile = 1.61 km, 3 Miles = 1.61 * 3 = 4.83 km

So, Nissan Leaf car can go 4.83 km/kWh

Hence, kWh it takes to go 1 km = 1/4.83 = 0.2 kWh

Anser 2.)

As per median estimates of US National Renewable Energy Laboratory for the emissions intensity of coal electricity is 0.97 kg CO2e/kWh.

Asswer 3.)

Cakculation for per km emmission from a standard American car:

Fuel economy of the car : 9.5 L/100 km or 0.095 L/km,

Life time carbon emmission empact of gasolene : 3.2 kg CO2e/L

Therefore, per km emmission from a standard American car shall be 0.095 * 3.2 = 0.304 kgCO2e